Out of a sample of 480 students, 310 indicated that they preferred chocolate ice cream to vanilla ice cream.

a. Estimate the value of the population proportion of those who preferred chocolate ice cream. (Round the final answer to 3 decimal places.)

Estimated population proportion           

b. Compute the standard error of the proportion. (Round the final answer to 4 decimal places.)

Standard error of the proportion           

c. Determine a 90% confidence interval for the population proportion of those who preferred chocolate ice cream. (Round the final answers to 3 decimal places.)

Confidence interval for the population proportion is between      and    .

d. Interpret your findings.    

If 480 such intervals were determined, the population  (Click to select)  mean  proportion  would be included in about  intervals.
Answer #1

Answer)

A)

P = 310/480 = 0.646

N = 480

B)

Standard error = √{p*(1-p)}/√n = 0.0218

C)

Critical value z from z table for 90% confidence level is 1.645

Margin of error (MOE) = Z*Standard error = 1.645*0.0218 = 0.035861

Interval is in between p-moe and p+moe

0.61 and 0.68

D)

We are 90% confident that true population proportion lies in this interval